OpenStudy (anonymous):

Is the period of y= cos x/4 going to be 2pi/4 or 2pi/1? ... I know the identity is 2pi/B... but i'm a little confused?... thanks

OpenStudy (anonymous):

Hm... \[\cos \frac{ x }{ 4 } = \cos x \frac{ 1 }{ 4 } = \frac{ 1 }{ 4 } \cos x\] You should be able to get the period once you switch everything around but... B now = 1/4 so I think you'll have to do it by hand if you don't want to get a decimal answer from your calculator? This is me guessing so I might be wrong.

OpenStudy (anonymous):

Oops, I put the 1/4 in front of cos instead of in front of x! My bad! Still the same thing as what I was talking about though ;(

OpenStudy (rational):

\(\large y = \cos(\frac{x}{4}) = \cos(\frac{1}{4}x) \) \(\large B = \frac{1}{4}\)

OpenStudy (rational):

like that @AltCode ? :)

OpenStudy (anonymous):

Yes that's what I meant! :)

OpenStudy (anonymous):

Ahhhh, ok. thank you so much. I see it. Great!! thanks

OpenStudy (anonymous):

thanks altcode

OpenStudy (rational):

then period will be \(\large \frac{2\pi}{\frac{1}{4}} = 8 \pi\)

OpenStudy (rational):

if u wana cheat : http://www.wolframalpha.com/input/?i=period+of+cos%28x%2F4%29 :P

OpenStudy (anonymous):

For \(\sin(x)\) we have: \[ 0\leq x<2\pi \]For \(\sin\left(\frac{x}4\right)\) we have:\[ 0\leq \frac x4 <2\pi \]

OpenStudy (anonymous):

I've tried using that wolfram alpha, but I can't seem to input the problems correctly on that site..

OpenStudy (rational):

we need to feed it in some specific format... :)

OpenStudy (anonymous):

Thanks. maybe someday I can help you guys out with something. WHO KNOWS!

OpenStudy (anonymous):

You don't get what I wrote...?

OpenStudy (anonymous):

Yes, i did. That made sense. It was clear when you set it up that way. I can use that each time. Thanks wio